What is the % w/V concentration of 65g of sugar (C6H12O6) dissolved in 750mL of water?
1
Expert's answer
2020-10-30T06:45:40-0400

solution:

Mass by Volume percentage (as solubilities) = (Mass of solute in grams/Volume of water in mL) x 100

Calculate w/v (%)

w/v (%) = (65.0 g ÷ 750 mL) × 100 = 8.67 g/100 mL (%) = 0.0867 g/mL (%)

Answer: w/v (%) = 8.67 g/100 mL (%)